I need help figuring out how to get "X" to one side. 2X/3=8+4X
Answer by edjones(8007)

2x/3=8+4x
2x=24+12x multiply each side by 3 to eliminate fraction.
-24=10x subtract 2x+24 from each side.
x=-24/10 divide each side by 10.
x=-2.4
.
Ed